]Question 4 (10 marks) Suppose the spot price of gold is $1,500 per troy ounce today. The futures price of gold for delivery i
Answer #1

Flat yield curve indicates no changes to interest rates

F=s(1+r)^n where f is futures rate

Spot price s = 1500

r is interest rate = we need to find

1530=1500(1+r)^1

1+r = 1.02

= 0.02 that is 2%

Theoretical futures price of 2 years

= 1500(1.02)^2 = 1560.6

B) after 3 months futures price is 1525

We shorted it with 10 contracts with 100troy per contract each

Present value of futures contract is = 1530/(1.02)^3/12 = 1522.444

Value of position is 10000+(1525-1522.444)×100×10

= 12556

Margin balance is 12556

Part c)

Futures price is 3000

Spot price = 2952

Risk free rate 2% dividend yield is 1% with annual compounding

F=S+i-d

I is interest and d is dividend yield

= 2952 + 2952(0.02)-2952(0.01)

= 2981.52

Actual futures price is more than theoretical price so futures is overpriced so arbitration exists and arbitration process is long index and short futures contract borrow money

In this strategy we will not loose money because we has hedged our position if index rises then gain is compensated by loss in futures and vice-versa and ultimately we gain 3000-2981.52= 18.48